### v1

# Kähler compactification of $\mathbb{C}^{n}$ and Reeb dynamics

Abstract.
Let $X$ be a smooth complex manifold. Assume that $Y\subset X$ is a Kähler submanifold such that $X\setminus Y$ is biholomorphic to $\mathbb{C}^{n}$. We prove that $(X,Y)$ is biholomorphic to $(\mathbb{P}^{n},\mathbb{P}^{n-1})$. We then study certain Kähler orbifold compactifications of $\mathbb{C}^{n}$ and, as an application, prove that on $\mathbb{C}^{3}$ the flat metric is the only asymptotically conical Ricci-flat Kähler metric whose metric cone at infinity has a smooth link. As a key technical ingredient, a new formula for minimal discrepancy of isolated Fano cone singularities in terms of Conley-Zehnder indices in symplectic geometry is derived.
1. Introduction
Our work is motivated by two problems in complex geometry. The first is Hirzebruch’s classical problem on analytic compactification of $\mathbb{C}^{n}$. Let $X$ be a compact complex manifold of complex dimension $n$. Assume that $Y$ is a subvariety of $X$ such that $X\setminus Y$ is biholomorphic to $\mathbb{C}^{n}$. Of course, the simplest example is $(X,Y)=(\mathbb{P}^{n},\mathbb{P}^{n-1})$ where $\mathbb{P}^{n-1}\subset\mathbb{P}^{n}$ is a linear subspace. In this case, we will just say that $(X,Y)$ is standard. In [Hir54], Hirzebruch asked the problem of classifying all such pair $(X,Y)$ with the second Betti number $1$. When $\dim X=1$ or $2$, it is known that there is only the standard example. For $n=3$, under the assumption that $X$ is projective, all such analytic compactifications of $\mathbb{C}^{3}$ are now classified. The subvariety $Y$ in any known non-standard example is always singular. For example, one can compactify $\mathbb{C}^{n}$ to a smooth quadric hypersurface in $\mathbb{P}^{n+1}$ by adding an irreducible singular divisor (a projective cone over a lower dimensional quadric) at infinity. It is a well-known conjecture that if $Y$ is assumed to be smooth then $(X,Y)$ should be standard (see [BM78], [PS91]). This was proved in [BM78] when $\dim X\leq 3$. Here we prove this conjecture under the Kähler assumption for all dimensions.
Theorem 1.1.
With the above notation, if $Y\subset X$ is a Kähler submanifold, $(X,Y)$ is standard.
Early results of Brenton-Morrow from [BM78] showed already that, with the same Kähler assumption, $Y$ must be a Fano manifold and they reduce the problem to proving that $Y$ is biholomorphic to $\mathbb{P}^{n-1}$. We will prove that the Fano index of $Y$ must be $n$ which then implies that $Y$ is $\mathbb{P}^{n-1}$ by a well-known criterion of Kobayashi-Ochiai ([KO73]). The main tools we use are ($S^{1}$-equivariant, positive) symplectic homology and two ways to compute it, one from the functorial properties and the special symplectic topology of $\mathbb{C}^{n}$, and the other from the Morse-Bott spectral sequences applied to the Boothy-Wang contact structure on the unit circle bundle of the normal bundle of $Y$.
Another problem that motivates our work is the classification of asymptotically conical (AC) Ricci-flat metrics, which has been recently studied extensively. Let $(W,g)$ be a complete Kähler manifold. It is called asymptotically conical (AC) if it is asymptotical to a Kähler cone $(\mathcal{C},g_{0})$ (see Definition 2.3). By the compactification results from [CH24, Li20] (see Theorem 1.3), the classification of such Kähler metrics are closely related to the orbifold compactifications of $W$. We have the following folklore conjecture  [^1] A version of such kind of conjecture was proposed by G. Tian. :
Conjecture 1.2.
Let $g$ be a complete AC Ricci-flat Kähler metric on $\mathbb{C}^{n}$ whose asymptotical cone $\mathcal{C}$ has a smooth link. Then $g$ is the flat Euclidean metric.
Note that by the recent examples constructed in [Li19, CR21, Sze19], the condition of having a smooth link is needed. We will use the symplectic techniques in an orbifold setting and reduce the conjecture 1.2 to Shokurov’s conjecture (see Conjecture 2.1) for isolated cone singularities.
Theorem 1.3.
The following statements are true:
  1. (1)
    Let $\mathcal{Y}\subset\mathcal{X}$ be an orbifold Kähler compactification of $\mathbb{C}^{n}$ that is associated to a complete AC Kähler metric. Then the orbifold cone $\mathcal{C}(\mathcal{Y},\mathcal{L})$ over $\mathcal{Y}$ with $\mathcal{L}$ being the orbifold normal bundle is a Gorenstein klt singularity which has the minimal discrepancy equal to $n-1$.
  2. (2)
    Conjecture 1.2 is true assuming Shokurov’s conjecture. In particular, Conjecture 1.2 is true when $n=3$.
As a key ingredient for proving the above results, we derive the following new formula for the minimal discrepancy of isolated Fano cone singularities (see 2.1 for the explanation of these terminologies) in terms of the $\mathrm{lSFT}$-invariant (see Definition 2.12), which was introduced by McLean ([McL16]) as a generalized Conley-Zehnder type index for closed Reeb orbits.
Theorem 1.4.
Let $o\in\mathcal{C}$ be an isolated Fano cone singularity of dimension $n$. For any quasi-regular conic contact form $\eta$ on the contact link $M$, we have the following formula for the minimal discrepancy:
(1) $22\mathrm{md}(o,\mathcal{C})=\inf_{\gamma}\mathrm{lSFT}_{\eta}(\gamma)>-$ 2 .
Here $\gamma$ on the right ranges over all closed Reeb orbits of $\eta$. If moreover $M$ admits a Liouville filling $W$ such that $c_{1}^{\mathbb{Q}}(W)=0\in H^{2}(W;\mathbb{Q})$, then we have
(2) $32\mathrm{md}(o,\mathcal{C})=\inf\{d\;|\;SH^{+,S^{1}}_{d}(W;\mathbb{Q})\neq 0\}+n-$ 3
where $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ denotes the $\mathbb{Q}$-coefficient $S^{1}$-equivariant positive symplectic homology of the Liouville filling $W$.
In the more special case of isolated quotient singularities, this formula recovers the result in [Zho24, Proposition 2.12]. We remark that our proof of Theorem 1.4 does not use the main formula from [McL16, Theorem 1.1] which states that $2\mathrm{md}$ is equal to the highest minimal index denoted by $\mathrm{hmi}$ for canonical singularities. In fact, the definition of the contact invariant $\mathrm{hmi}$ in [McL16] requires enumeration through all contact forms, while 1.4 asserts that the minimal discrepancy can be computed using the Reeb dynamics from a single conic contact form and is witnessed by a non-trivial class in Floer homology. The latter can be approached using properties of symplectic homology, which allow us to show the equality $\mathrm{md}(o,\mathcal{C})=n-1$ for the Fano cone arising in our problems. The existence of a Liouville filling $W$ is just for the existence of $\mathbb{Q}$-graded $S^{1}$-equivariant positive symplectic cohomology. Such a Liouville filling always exists if $o$ is smoothable and $\dim_{\mathbb{C}}\mathcal{C}\geq 3$. The specific choice of $W$ does not matter. In a more general setting, one can replace $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ by the linearized contact homology of $M$ w.r.t. a $\mathbb{Q}$-graded augmentation over $\mathbb{C}$, which always exists in this case, see 3.8 for more discussions.
The application of tools from symplectic geometry to compactification problem seems new to us. From a broader perspective, we use techniques from symplectic geometry to rule out certain singularities. The application of (very different) symplectic techniques to complex geometry to rule out singularities has appeared in the recent solution to the generalized Bishop problem [FHYZ24].
Acknowledgments
C. Li is partially supported by NSF (Grant No. DMS-2305296). Z. Zhou is supported by National Key R&D Program of China under Grant No.2023YFA1010500, NSFC-12288201 and NSFC-12231010. C. Li would like to thank Prof. Laszlo Lempert for helpful discussions, and Prof. Mircea Mustaţă for comments on Shokurov’s conjecture. Both authors would like to thank Prof. Xiaojun Huang for putting us into connection.
2. Preliminaries
2.1. Fano cone singularities and AC Kähler metrics
Let $\mathcal{C}$ be a normal affine variety with log terminal $\mathbb{Q}$-Gorenstein singularities. Here $\mathbb{Q}$-Gorenstein means that the canonical divisor $K_{\mathcal{C}}$ is $\mathbb{Q}$-Cartier. To define the log terminal condition, we choose a log resolution of singularities $\mu:\tilde{\mathcal{C}}\rightarrow\mathcal{C}$ such that the inverse image of the singular locus of $\mathcal{C}$ is given by a simple normal crossing divisor $\bigcup_{i}E_{i}$. We can then write an identity:
$K_{\tilde{\mathcal{C}}}=\mu^{*}K_{\mathcal{C}}+\sum_{i}a(E_{i},\mathcal{C})E_{i}.$
The affine variety $\mathcal{C}$ has log terminal singularities if $a(E_{i},\mathcal{C})>-1$ for any $i$. Log terminal singularities are local analogues of Fano varieties and they play an important role in birational algebraic geometry. From now on, we always assume that $o\in\mathcal{C}$ is an isolated singularity. So we can choose a resolution of singularity $\mu:\tilde{\mathcal{C}}\rightarrow\mathcal{C}$ that is an isomorphism in a punctured neighborhood of $o\in\mathcal{C}$ such that $\mu^{-1}(o)$ is a simple normal crossing divisor. The minimal discrepancy of the klt singularity $o\in\mathcal{C}$ is then defined as $\mathrm{md}(o,\mathcal{C})=\min_{i}\{a(E_{i},\mathcal{C});\mu(E_{i})=o\}$. One can show that this quantity does not depend on the choice of log resolutions. We have the following important conjecture:
Conjecture 2.1 (Shokurov).
For any klt singularity $o\in\mathcal{C}$ of dimension $n$, $\mathrm{md}(o,\mathcal{C})\leq n-1$ and the equality holds if and only if $o\in\mathcal{C}$ is a smooth point.
Shokurov’s conjecture has been proved in dimension $\leq 3$ ([Mar96, Kaw93]).
We now explain the concept of a Kähler cone metric on $\mathcal{C}$. This means that there is smooth function $\mathfrak{r}:\mathcal{C}^{\circ}:=\mathcal{C}\setminus o\rightarrow\mathbb{R}_{>0}$, called the radius function, which is a surjective submersion such that if we set $\hat{\omega}:=\sqrt{-1}\partial\bar{\partial}{\mathfrak{r}}^{2}$ and $g_{0}=\hat{\omega}(\cdot,J_{0}\cdot)$ where $J_{0}$ is the integrable complex structure on $\mathcal{C}^{\circ}$, then $g_{0}$ is a Riemannian metric on $\mathcal{C}^{\circ}$ that is isometric to a Riemannian cone metric $d{\mathfrak{r}}^{2}+{\mathfrak{r}}^{2}g^{M}$ where $M=\{{\mathfrak{r}}=1\}$ is the link and $g^{M}=g_{0}|_{M}$. Define the conic contact form associated to the radius function ${\mathfrak{r}}$ as $\eta=-{\mathfrak{r}}^{-1}J_{0}d{\mathfrak{r}}$. Then we have $\hat{\omega}=2{\mathfrak{r}}d{\mathfrak{r}}\wedge\eta+{\mathfrak{r}}^{2}d\eta$. The Reeb vector field, denoted by ${\mathfrak{v}}=J_{0}(\mathfrak{r}\partial_{\mathfrak{r}})$, is a holomorphic Killing vector field on $\mathcal{C}$ and satisfies $\eta({\mathfrak{v}})=1$. The associated holomorphic vector field ${\mathfrak{v}}_{\mathbb{C}}={\mathfrak{r}}\partial_{\mathfrak{r}}-\sqrt{-1}J_{0}({\mathfrak{r}}\partial_{\mathfrak{r}})$ generates an effective holomorphic action by a torus $(\mathbb{C}^{*})^{m}$ which we will also denote by $\langle{\mathfrak{v}}\rangle$. If $m=1$, ${\mathfrak{v}}$ and $\eta$ are called quasi-regular and otherwise (i.e. when $m\geq 2$) ${\mathfrak{v}}$ and $\eta$ are called irregular.
Remark 2.2.
As the notation conventions in the Sasaki geometry literature and contact geometry literature are different, we fix $\eta$ as the contact form, ${\mathfrak{v}}$ as the Reeb vector field, and $\xi:=\ker\eta$ as the contact structure in this paper.
In the quasi-regular case, we can take the quotient of $\mathcal{C}^{\circ}$ by the $\mathbb{C}^{*}$-action or equivalently the quotient of $M$ by the $S^{1}$-action to get an orbifold $\mathcal{Y}:=(Y,\Delta)=M/S^{1}=\mathcal{C}^{\circ}/\mathbb{C}^{*}$ equipped with an orbifold line bundle $\mathcal{L}$. There is a natural projection $\pi:\mathcal{C}^{\circ}\rightarrow Y$ which makes $\mathcal{C}^{\circ}$ a $\mathbb{C}^{*}$-Seifert bundle over the cyclic orbifold $\mathcal{Y}$ which was studied for example by Kollár in [Kol04]. The affine variety $\mathcal{C}$ is algebraically described as:
$\mathcal{C}=\mathcal{C}(\mathcal{Y},\mathcal{L}):=\mathrm{Spec}\left(\bigoplus_{m}H^{0}(\mathcal{Y},m\mathcal{L})\right).$
Geometrically it is obtained from the total space of the orbifold line bundle $\mathcal{L}^{-1}$ by contracting the zero section. For simplicity, We will call $o\in\mathcal{C}$ an orbifold cone singularity. $\mathcal{C}$ can be compactified to a projective variety $\overline{\mathcal{C}}$ by adding an divisor $\mathcal{Y}_{\infty}\cong\mathcal{Y}$ at infinity (see [Kol04, 14]). According to [Kol04, 42], the singularity $o\in\mathcal{C}$ is log terminal if and only $(Y,\Delta)$ is a Fano orbifold (i.e. the orbifold anti-canonical line bundle $-K^{\mathrm{orb}}_{(Y,\Delta)}$ is ample) and there exists $\ell\in\mathbb{Q}_{>0}$ such that $-K^{\mathrm{orb}}_{(Y,\Delta)}\sim_{\mathbb{Q}}\ell\mathcal{L}$ (see Proposition 2.10 for a proof in our setting). In this case, we will simply call the orbifold cone $o\in\mathcal{C}$ a Fano cone (following the terminology in [LLX20, Definition 3.7]). Moreover, in the quasi-regular case, the metrics $g_{0}$ and $g^{M}$ discussed in the previous paragraph are induced by a Hermitian metric $h$ on $\mathcal{L}$ over $(Y,\Delta)$ whose Chern curvature will be denoted by $\omega^{Y}$.
We now introduce the class of Kähler metrics on quasi-projective varieties that are modeled on Kähler cone metrics near infinity.
Definition 2.3.
Let $(W,g)$ be a complete Kähler manifold whose integrable complex structure is denoted by $J$. It is called asymptotically conical (AC) with the asymptotical cone $(\mathcal{C},g_{0})$ if there exists a compact subset $K\subset W$ and a diffeomorphism $\Phi:\{{\mathfrak{r}}>1\}\rightarrow M\setminus K$ such that for some $\lambda_{1}>0,\lambda_{2}>0$ and all $j\in\mathbb{Z}_{\geq 0}$.
(3) $|\nabla^{j}_{g_{0}}(\Phi^{*}g-g_{0})|_{g_{0}}=O({\mathfrak{r}}^{-\lambda_{1}-j}),$
(4) $|\nabla^{j}_{g_{0}}(\Phi^{*}J-J_{0})|_{g_{0}}=O({\mathfrak{r}}^{-\lambda_{2}-j}).$
If the Kähler metric $g$ is Ricci-flat, then its asymptotical cone $(\mathcal{C},g_{0})$ is also Ricci-flat. In this case, the metric $g^{M}$ on the link $M$ is called Sasaki-Einstein. When ${\mathfrak{v}}$ is quasi-regular, then $g^{M}$ is Sasaki-Einstein if and only if the orbifold Kähler metric $\omega^{Y}$ is Kähler-Einstein i.e. $Ric(\omega^{Y})=\ell\omega^{Y}$. We refer to the book [BG08] for these facts and an extensive exposition on Sasaki geometry, particularly Sasaki-Einstein metrics.
We will use the following compactification result which only assumes that the complex structure is asymptotical to the cone.
Theorem 2.4.
Let $(W,g,J)$ be a complete Kähler manifolds and $(\mathcal{C},g_{0},J_{0})$ be an orbifold cone as before. Assume that there exist a compact subset $K\subset W$ and a diffeomorphism $\Phi:\{\mathfrak{r}>1\}\rightarrow M\setminus K$ such that for some $\lambda_{2}>0$ and all $j\in\mathbb{Z}_{\geq 0}$,
$|\nabla^{j}_{g_{0}}(\Phi^{*}J-J_{0})|_{g_{0}}=O(\mathfrak{r}^{-\lambda_{2}-j}).$
  1. (1)
    ([Li20, CH24]) Assume that the Reeb vector field ${\mathfrak{v}}$ is quasi-regular on the cone $\mathcal{C}$. There exists a compact Kähler orbifold $\mathcal{X}$ and a sub-orbifold divisor $\mathcal{Y}$ such that $\mathcal{Y}\cong\mathcal{C}^{\circ}/\langle\mathfrak{v}\rangle$ and $\mathcal{X}\setminus\mathcal{Y}$ is biholomorphic to $W$. Moreover there exists an orbifold diffeomorphism $\Phi:(\mathcal{U},\mathcal{Y})\rightarrow(\mathcal{V},\mathcal{Y}_{\infty})$ where $\mathcal{U}$ is a neighborhood of $\mathcal{Y}\subset\mathcal{X}$ and $\mathcal{V}$ is a neighborhood of $\mathcal{Y}_{\infty}\subset\overline{\mathcal{C}}$.
  2. (2)
    ([CH24]) Assume that the Reeb vector field ${\mathfrak{v}}$ is irregular. Then there exists a sequence of radius functions $\mathfrak{r}_{i}$ on $\mathcal{C}$ whose associated Reeb vector fields ${\mathfrak{v}}_{i}$ are quasi-regular such that ${\mathfrak{v}}_{i}$ converge to ${\mathfrak{v}}$ as $i\rightarrow+\infty$. For any $0<\epsilon\ll\lambda_{2}$ there exists $I$ such that for any $i\geq I$, $|\nabla_{g_{0,i}}^{j}(\Phi^{*}J-J_{0})|=O(\mathfrak{r}_{i}^{-\lambda_{2}+\epsilon-j})$ for all $j\in\mathbb{Z}_{\geq 0}$. As a consequence, there exists orbifold compactification of $W$ by $\mathcal{Y}_{i}=\mathcal{C}^{\circ}/\langle{\mathfrak{v}}_{i}\rangle$ when $i\geq I$.
[…]
2.2. Orbifold Kähler compactification of $\mathbb{C}^{n}$
We generalize some of Brenton-Morrow’s results to the orbifold setting. Let $(\mathbb{C}^{n},g)$ be an AC Kähler metric that is asymptotical to $(\mathcal{C},g_{0})$ with a quasi-regular Reeb vector field ${\mathfrak{v}}$. Let $\mathcal{Y}\subset\mathcal{X}$ be the orbifold compactification of $\mathbb{C}^{n}$ obtained by Theorem 2.4. In particular, $\mathcal{Y}=\mathcal{C}/\langle{\mathfrak{v}}\rangle$. Let $[\mathcal{Y}]$ denote the orbifold line bundle on $\mathcal{X}$ associated to the orbifold divisor $[\mathcal{Y}]$ and denote by $\mathcal{L}$ its restriction to $\mathcal{Y}$ which coincides with the orbifold normal bundle $\mathcal{L}=[\mathcal{Y}]|_{\mathcal{Y}}=N_{\mathcal{Y}/\mathcal{X}}$. We will use the orbifold cohomology introduced in [Hae84]. In other words, we set $H^{i}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})=H^{i}(B\mathcal{X},\mathbb{Z})$ and $H^{i}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=H^{i}_{\mathrm{orb}}(B\mathcal{Y},\mathbb{Z})$ where $B\mathcal{X}$ and $B\mathcal{Y}$ are the classifying spaces of the orbifolds $\mathcal{X}$ and $\mathcal{Y}$ respectively. $B\mathcal{X}$ can be constructed by the following procedure (see [Hae84, HS91] and [BG08, Chapter 4] for more details). Choose a Riemannian metric on $\mathcal{X}$ and let $P$ denote the differentiable bundle of orthonormal frames of $\mathcal{X}$. Let $EO(2n)\rightarrow BO(2n)$ be a universal $O(2n)$-bundle, then $B\mathcal{X}$ is the quotient of $EO(2n)\times P$ by the diagonal action of $O(2n)$. A similar construction applies to $\mathcal{Y}$. Moreover, since it is a global quotient of $M$ by an effective $S^{1}$ action we can use the following construction following [HS91, section 5.3]. We choose an $S^{1}$-invariant Riemannian metric on $M$. Let $P_{M}$ denote the bundle of orthonormal frames of $M$ and set $BM=P_{M}\times_{O(2n-1)}EO(2n-1)$ where $EO(2n-1)\rightarrow BO(2n-1)$ is a universal $O(2n-1)$-bundle. We can set $B\mathcal{Y}=BM/S^{1}$ and we have a principal $S^{1}$-bundle $BM\rightarrow BM/S^{1}=B\mathcal{Y}$.
Proposition 2.5.
With the above notation, the following properties hold true:
  1. (1)
    There are ring isomorphisms $H_{\mathrm{orb}}^{*}(\mathcal{X},\mathbb{Q})\cong H^{*}(\mathbb{P}^{n},\mathbb{Q})$ and $H_{\mathrm{orb}}^{*}(\mathcal{Y},\mathbb{Q})\cong H^{*}(\mathbb{P}^{n-1},\mathbb{Q})$.
  2. (2)
    $H^{2}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})$ is generated by $c_{1}^{\mathrm{orb}}([\mathcal{Y}])$ and $H^{2}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})$ is generated by $c_{1}^{\mathrm{orb}}(\mathcal{L})$.
  3. (3)
    $-K^{\mathrm{orb}}_{\mathcal{X}}=(r+1)[\mathcal{Y}]$ with $r\geq 1$ is an ample orbifold line bundle and $-K^{\mathrm{orb}}_{\mathcal{Y}}=r\mathcal{L}$ is also ample. In particular, both $\mathcal{X}$ and $\mathcal{Y}$ are Fano orbifolds and $(r+1)$, $r$ are their Fano indices respectively.
  4. (4)
    The link $M$ of the cone $\mathcal{C}$ is an integral homology sphere and bounds a homology ball in $\mathbb{C}^{n}$.
Proof.
Let $\overline{\mathcal{U}}$ be the closed neighborhood of $\mathcal{Y}\subset\mathcal{X}$ as obtained from Theorem 2.4. We denote by $\overline{U}$ the underlying space of orbifold $\overline{\mathcal{U}}$ (with smooth boundary). We have the exact sequence (for ordinary cohomology):
(5) $H^{k}(X-U,\partial\overline{U},\mathbb{Z})\rightarrow H^{k}(X-\overline{U},\mathbb{Z})\rightarrow H^{k}(\partial\overline{U},\mathbb{Z})\rightarrow H^{k+1}(X-\overline{U},\partial\overline{U},\mathbb{Z}).$
Using excision property, $H^{k}(X-\overline{U},\partial\overline{U},\mathbb{Z})=H^{k}(X,\overline{U},\mathbb{Z})=H^{k}(X,Y,\mathbb{Z})=H^{k}_{c}(X\setminus Y,\mathbb{Z})$ for $0\leq k<2n$. So we have
(6) $H^{k}(X-\overline{U},\mathbb{Z})\cong H^{k}(\partial\overline{U},\mathbb{Z})\quad\text{ for }0\leq k\leq 2n-2.$
On the other hand, by Lefschetz duality, $H_{k}(X-\overline{U},\mathbb{Z})=H^{2n-k}(X-\overline{U},\partial\overline{U},\mathbb{Z})=0$ for $1\leq k\leq 2n$. By Universal Coefficients Theorem, $H^{k}(X-\overline{U},\mathbb{Z})=0$ for $1\leq k\leq 2n$. So we also get:
(7) $H^{k}(\partial\overline{U},\mathbb{Z})=0\quad\text{for}\quad 1\leq k\leq 2n-2.$
Next, we consider the exact sequence for relative orbifold cohomology:
(8) $H^{k}_{\mathrm{orb}}(\mathcal{X},\mathcal{Y},\mathbb{Z})\rightarrow H^{k}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\rightarrow H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})\rightarrow H^{k+1}_{\mathrm{orb}}(\mathcal{X},\mathcal{Y},\mathbb{Z})$
We have $H^{k}_{\mathrm{orb}}(\mathcal{X},\mathcal{Y},\mathbb{Z})=H^{k}_{c}(\mathcal{X}\setminus\mathcal{Y},\mathbb{Z})=0$ for $0\leq k<2n$. So $H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=H^{k}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})$ for $20\leq k\leq 2n-$ 2. In particular $H^{0}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=H^{0}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\cong\mathbb{Z}$ and hence $Y$ is connected. Since $H^{2n}_{\mathrm{orb}}(\mathcal{X},\mathcal{Y},\mathbb{Z})=H^{2n}_{c}(X\setminus Y,\mathbb{Z})=\mathbb{Z}$, when $k=2n-1$, we get:
(9) $0\rightarrow H^{2n-1}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\rightarrow H^{2n-1}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})\rightarrow\mathbb{Z}{\rightarrow}H^{2n}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\rightarrow H^{2n}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})\rightarrow 0$
For $k\geq 2n+1$, we get $H^{k}_{\mathrm{orb}}(\mathcal{X})\cong H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})$.
Let $M$ denote the circle bundle of $\mathcal{L}$, which is diffeomorphic to $\partial\overline{U}$. We apply the Leray spectral sequence to the $S^{1}$-principal bundle $BM\rightarrow B\mathcal{Y}$ to get the Gysin exact sequence:
(10) $H^{k}(M,\mathbb{Z})\rightarrow H^{k-1}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})\stackrel{{\scriptstyle\alpha_{k-1}}}{{\longrightarrow}}H^{k+1}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})\rightarrow H^{k+1}(M,\mathbb{Z})$
where $\alpha_{k-1}$ is the cup product with $c_{1}(\mathcal{L})\in H^{2}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})$. By the vanishing property (7), we know that $\alpha_{k-1}$ is an isomorphism for $31\leq k\leq 2n-$ 3. Since $H^{0}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=\mathbb{Z}$, we get $H^{p}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=c_{1}^{\mathrm{orb}}(\mathcal{L})^{p/2}\mathbb{Z}$ for $20\leq p\leq 2n-$ 2 even. For $k=0$ in (10), we get $H^{1}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=0$ and we get $H^{p}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=0$ when $31\leq p\leq 2n-$ 3 is odd. By applying the same argument to $\mathbb{Q}$ coefficients, we also see that the orbifold cohomology ring of $\mathcal{Y}$ with $\mathbb{Q}$-coefficients is the same as the cohomology ring of $\mathbb{P}^{n-1}$ which is generated by $c_{1}(\mathcal{L})$.
Next, the Kähler assumption implies that there is a Hodge decomposition of $H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{C})$ and $H^{k}_{\mathrm{orb}}(\mathcal{X},\mathbb{C})$ for any $k$ (see [PS08, Part I, 2.5]). In particular, we have
(11) $H^{k}_{\mathrm{orb}}(\mathcal{X},\mathbb{C})=\bigoplus_{p+q=k}H^{p}_{\mathrm{orb}}(\mathcal{X},\Omega^{q}_{\mathcal{X}}).$
We get in particular that $H^{1}_{\mathrm{orb}}(\mathcal{X},\mathcal{O}_{\mathcal{X}})=0$ and $H^{2}_{\mathrm{orb}}(\mathcal{X},\mathcal{O}_{X})=0$. The exponential sequence in the complex orbifold setting (see [BG08, Theorem 4.4.23]) implies that the map $\mathrm{Pic}^{\mathrm{orb}}(\mathcal{X})=H^{1}_{\mathrm{orb}}(\mathcal{X},\mathcal{O}^{*}_{\mathcal{X}})\cong H^{2}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\cong H^{2}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=\mathbb{Z}$. Because $H^{1}(M;\mathbb{Z})=0$, we know that $c_{1}(\mathcal{L})$ is primitive in $H^{2}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})$.
This implies in particular, $-K^{\mathrm{orb}}_{\mathcal{X}}=b[\mathcal{Y}]$ for some $b\in\mathbb{Z}$. If $b\leq 0$, then $K^{\mathrm{orb}}_{\mathcal{X}}$ has nonzero holomorphic section $s_{\mathcal{Y}}^{-b}$ which contradicts $\dim H^{0}_{\mathrm{orb}}(\mathcal{X},K^{\mathrm{orb}}_{\mathcal{X}})=h^{n,0}_{\mathrm{orb}}(X)=0$. So $b=1+r\geq 1$ i.e. $r\geq 0$. By adjunction formula, we get $-K^{\mathrm{orb}}_{\mathcal{Y}}=r\mathcal{L}$. If $r=0$, then $K^{\mathrm{orb}}_{\mathcal{Y}}$ is trivial and $H^{n-1}_{\mathrm{orb}}(\mathcal{Y},\mathcal{O}_{Y})=H^{0}_{\mathrm{orb}}(\mathcal{Y},K^{\mathrm{orb}}_{\mathcal{Y}})\cong\mathbb{C}$ which contradicts to the vanishing $h^{0,n-1}(\mathcal{Y})=0$. So $r\geq 1$.
By (7), we see that $M=\partial\bar{U}$ is an integral homology sphere. By the Mayer-Vietoris sequence to $\mathbb{C}^{n}=\overline{U}\cup(\mathbb{C}^{n}\setminus\overline{U})$, we know that $M=\overline{U}\cap(\mathbb{C}^{n}\setminus\overline{U})$ bounds an integral homology ball. ∎
Remark 2.6.
In general $H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})$ is not 0 when $k\geq 2n-1$. For example, it is well-known that for the weighted projective space $\mathbb{P}({\bf w})=\mathbb{P}(w_{1},w_{2},\dots,w_{n})$ with $d=\prod_{i}w_{i}$, we have the orbifold cohomology:
(12) $H^{k}_{\mathrm{orb}}(\mathbb{P}({\bf w}),\mathbb{Z})=\left\{\begin{array}[]{ll}\mathbb{Z}&k\leq 2n-2\text{ and is even}\\ \mathbb{Z}_{d}&k\geq 2n\text{ and is even}\\ 0&k\text{ is odd}.\end{array}\right.$
Definition 2.7.
Given two contact manifolds $(M_{+},\xi_{+}),(M_{-},\xi_{-})$, we say $(W,\lambda)$ is a Liouville cobordism from $(M_{-},\xi_{-})$ to $(M_{+},\xi_{+})$ if
  1. (1)
    $\mathrm{d}\lambda$ is a symplectic form on $W$;
  2. (2)
    $\partial W=(-M_{-})\cup M_{+}$ as oriented manifolds;
  3. (3)
    The Liouville vector field $X_{\lambda}$ defined by $\lambda=\iota_{X_{\lambda}}\mathrm{d}\lambda$ is transverse to $\partial W$ pointing outward/inward along $M_{+}/M_{-}$ respectively;
  4. (4)
    $\ker\lambda|_{M_{\pm}}=\xi_{\pm}$.
$(W,\lambda)$ is called a Liouville filling of $(M_{+}.\xi_{+})$ if $M_{-}=\emptyset$.
If $M$ is the smooth boundary of strongly pseudo-convex domain $W$, then $(W,-\mathrm{d}^{\mathbb{C}}\rho)$ is a Liouville filling of $M$ whose contact structure is the CR structure, where $\rho$ is the strongly plurisubharmonic function whose sublevel let is $W$ with $M$ a regular level set. More generally, $(\rho^{-1}([a,b]),-\mathrm{d}^{\mathbb{C}}\rho)$ is a Liouville cobordism from $\rho^{-1}(a)$ to $\rho^{-1}(b)$ assuming both $a,b$ are regular values. In particular, by [H7̈3, Theorem 2.6.12], we have the following.
Proposition 2.8.
The domain $W$ bounded by the contact link $M$ in $\mathbb{C}^{n}$ by 2.5 gives a Liouville filling of $M$.
In general, being a strongly pseudo-convex hypersurface in $\mathbb{C}^{n}$ does not imply the cobordism from $M$ to the standard $S^{2n-1}$ end of $\mathbb{C}^{n}$ is a Liouville cobordism. Thanks to the asymptotic cone structure here, we do have a Liouville cobordism in our case, which is crucial for our computation of symplectic homology.
Proposition 2.9.
The cobordism from the contact link $M$ to the standard $S^{2n-1}$ end of $\mathbb{C}^{n}$ is a Liouville cobordism from the CR structure on $M$ to the standard contact/CR structure on $S^{2n-1}$.
Proof.
Since $M$, as a contact manifold, can be viewed as a level set of the radius function $\mathfrak{r}^{2}$, where $\mathfrak{r}^{2}$ defines a strongly plurisubharmonic function on the end of $\mathbb{C}^{n}$ by the asymptotic cone assumption. Let $\rho=|z|^{2}$ using the Euclidean metric, then $f(\rho)$ is plurisubharmonic on $\mathbb{C}^{n}$ if $f^{\prime\prime}\geq 0,f^{\prime}\geq 0$. We assume $M$ is contained in the ball $B(r)$ for $r\gg 0$. We can choose $f$, such that $f^{\prime\prime}\geq 0,f^{\prime}\geq 0$, $f(\rho)=0$ for $\rho\leq r^{2}$. Moreover, we require $f$ such that there exists $R\gg r$ so that $|\nabla\mathfrak{r}^{2}|\ll|\nabla f(\rho)|$ outside the ball $B(R)$. As a consequence, for $C\gg 0$, $C$ is a regular value of $F_{t}:=t\mathfrak{r}^{2}+f(\rho)$ for all $t\in[0,1]$. $F_{1}=\mathfrak{r}^{2}+f(\rho)$ is strongly plurisubharmonic outside $M$ and inside $F_{1}^{-1}(C)$, where $M$ is a regular level set of $F_{1}$ by construction. Therefore we have a Liouville cobordism from $M$ to $F_{1}^{-1}(C)$. Finally, as $F_{t}^{-1}(C)$ is a smooth family of strongly pseudo-convex hypersurfaces in $\mathbb{C}^{n}$, the Gray stability implies that $F_{1}^{-1}(C)=F_{0}^{-1}(C)$ as contact manifolds, while $F_{0}^{-1}(C)$ is the standard contact sphere by construction. ∎
Proposition 2.10.
The vertex $o\in\mathcal{C}$ is a Gorenstein klt singularity.
Proof.
Our argument is similar to [Kol13, 3.1]. Let $\mathcal{C}^{\prime}$ denote the total space of the orbifold line bundle $\pi_{\mathcal{C}^{\prime}}:\mathcal{L}^{-1}\rightarrow\mathcal{Y}$. Then we have the identity $K_{\mathcal{C}^{\prime}}\otimes\mathcal{O}(\mathcal{Y}_{0})=\pi_{\mathcal{C}^{\prime}}^{*}K^{\mathrm{orb}}_{\mathcal{Y}}$. Away from the zero section $\mathcal{Y}_{0}$ we have $K_{\mathcal{C}^{\circ}}=\pi^{*}K^{\mathrm{orb}}_{\mathcal{Y}}=\pi^{*}(\mathcal{L}^{-r})\cong\mathcal{O}_{\mathcal{C}^{\circ}}$ where $\pi:\mathcal{C}^{\circ}\rightarrow\mathcal{Y}$ is the natural projection. Since $\mathcal{C}$ is normal, $K_{\mathcal{C}}$ is also trivial and hence $\mathcal{C}$ is Gorenstein.
Let $\mu_{1}:\mathcal{C}^{\prime}\rightarrow\mathcal{C}$ be the contraction of the zero section. We have the identity:
$K_{\mathcal{C}^{\prime}}=\mu_{1}^{*}K_{\mathcal{C}}+(r-1)\mathcal{Y}_{0}$
where $\mathcal{Y}_{0}$ is the zero section of $\mathcal{L}^{-1}$. Let $\mu_{2}:\tilde{\mathcal{C}}\rightarrow\mathcal{C}^{\prime}$ be the resolution of singularities with simple normal crossing exceptional divisors $\{F_{i}\}$, we get:
$K_{\tilde{\mathcal{C}}}+\widetilde{\mathcal{Y}}_{0}$ $=$ $\mu_{2}^{*}(K_{\mathcal{C}^{\prime}}+\mathcal{Y}_{0})+\sum_{i}a_{i}F_{i}$
$=$ $\mu_{2}^{*}\mu_{1}^{*}K_{\mathcal{C}}+r\mu_{2}^{*}\mathcal{Y}_{0}+\sum_{i}a_{i}F_{i}$
$=$ $\mu^{*}K_{\mathcal{C}}+r\widetilde{\mathcal{Y}}_{0}+\sum_{i}(a_{i}+r\cdot\mathrm{ord}_{F_{i}}(\mathcal{O}(-\mathcal{Y}_{0}))F_{i},$
where $\mathcal{O}(-\mathcal{Y}_{0})$ denotes the ideal sheaf of $\mathcal{Y}_{0}\subset\mathcal{C}^{\prime}$ and $\widetilde{\mathcal{Y}}_{0}$ is the strict transform of $\mathcal{Y}_{0}$ under $\mu_{2}$ and $\mu=\mu_{2}\circ\mu_{1}:\tilde{\mathcal{C}}\rightarrow\mathcal{C}$ is a resolution of singularity. Note that $\mathcal{Y}_{0}$ has only quotient singularities which are klt. By inversion of adjunction (see [Kol04, Theorem 4.9] and the references there), $(\mathcal{C}^{\prime},\mathcal{Y}_{0})$ is plt which implies $a_{i}=a(F_{i},\mathcal{C}^{\prime},\mathcal{Y}_{0})>-1$. We see that $o\in\mathcal{C}$ is klt since the coefficients of the exceptional divisors of $\mu$ are bigger than $-1$. ∎
We can also write the minimal discrepancy of $o\in\mathcal{C}$ as
(13) $\text{md}(o\in\mathcal{C})=\min\{r-1,a(F_{i},\mathcal{C}^{\prime},(1-r)\mathcal{Y}_{0})\}.$
We can then derive an algebraic formula for the minimal discrepancy of $o\in\mathcal{C}$ by adapting the argument as in [Kol13, Proof of Theorem 3.21] as follows.
First note that $\mathcal{C}^{\prime}$ has only quotient singularities that are contained in $\mathcal{Y}_{0}$. Let $p\in\mathcal{C}^{\prime}$ be a quotient singularity given by $\pi_{p}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}/G_{p}$ where $G_{p}$ is a cyclic subgroup of $U(n)$. According to the local classification of smooth Seifert $\mathbb{C}^{*}$-bundles from [Kol04, 25], $\mathbb{C}^{n}/G_{p}$ is isomorphic to $\mathbb{C}^{n}/\frac{1}{m}(b_{1},\dots,b_{n})$ that satisfies the condition that $\pi_{p}^{*}\mathcal{Y}_{0}=\{x_{1}=0\}$ with $\{x_{1},\dots,x_{n}\}$ being linear coordinates on $\mathbb{C}^{n}$ and $b_{1}$ is relatively prime to $m$. We have the following formula for the minimal discrepancy of $o\in\mathcal{C}$ which can be seen as a generalization of Reid-Tai’s criterion for klt quotient singularities to the case of isolated Fano cone singularities.
Proposition 2.11.
With the above notations, we have the equality
(14) $\mathrm{md}(o,\mathcal{C})=\min_{p,g}\left\{r,\frac{1}{m}\left(\sum_{i=1}^{n}w_{i}(g)+(r-1)w_{1}(g)\right)\right\}-1$
where $p$ ranges over all quotient singularities on $\mathcal{C}^{\prime}$ and $g$ ranges over all non-identity elements in the cyclic groups $G_{p}\cong\mathbb{Z}_{m}$ and satisfies $g^{*}x_{i}=\epsilon^{w_{i}(g)}x_{i}$ with $\epsilon=e^{2\pi\sqrt{-1}/m}$ and $0\leq w_{i}(g)<m$.
Proof.
We pick any exceptional divisor $F$ over $\mathbb{C}^{n}/G$ dominated by an exceptional divisor $E$ over $\mathbb{C}^{n}$ that is pointwise fixed by a cyclic subgroup $\langle g\rangle$ of order $k>1$. Assume $m=k\ell$. We have the equality $a(F,\mathbb{C}^{n}/G,(1-r)\mathcal{Y}_{0})+1=k^{-1}(a(E,\mathbb{C}^{n},\Delta^{\prime})+1)$ where $\Delta^{\prime}=(1-r)\pi_{p}^{*}\mathcal{Y}_{0}$. At the generic point of $E$ choose local coordinates $y_{1},\dots,y_{n}$ such that $E=\{y_{1}=0\}$, $g^{*}y_{1}=\epsilon^{\ell}y_{1}$ is a $g$-eigenfunction (up to a change of the generator of $\langle g\rangle$). Using the linear coordinates $\{x_{i}\}$ on $\mathbb{C}^{n}$ as above such that $\pi_{p}^{*}\mathcal{Y}_{0}=\{x_{1}=0\}$, we can write $x_{i}=y_{1}^{c_{i}}u_{i}$ where the $u_{i}$ are units. Then $g^{*}(dx_{1}\wedge\cdots\wedge dx_{n})$ vanishes to order $\geq(-1+\sum_{i}c_{i})$. If $g^{*}x_{i}=\epsilon^{w_{i}}x_{i}$ with $0\leq w_{i}<m$, then $\ell c_{i}\equiv w_{i}\text{ mod }m$. So we get:
$a(F,\mathbb{C}^{n}/G,(1-r)\mathcal{Y}_{0})+1$ $=$ $\frac{1}{k}(a(E,\mathbb{C}^{n},(1-r)\mathcal{Y}_{0})+1)$
$\geq$ $\frac{1}{k}(\sum_{i}c_{i}+(r-1)c_{1})\geq\frac{1}{k\ell}(\sum_{i}w_{i}+(r-1)w_{1}).$
Because of (13), this proves that the left-hand-side of (14) is greater than the right-hand-side.
Conversely any element $g\in G$ generates a cyclic subgroup which acts faithfully on $\mathbb{C}^{n}$ as $\mathbb{C}^{n}/\frac{1}{m}(w_{1},\dots,w_{n})$ where $\mathcal{Y}_{0}=\{x_{1}=0\}$. We want to show that $\frac{1}{m}(\sum_{i}w_{i}+(r-1)w_{1})\geq\mathrm{md}(o,\mathcal{C})$. If $\mathrm{gcd}(w_{1},\dots,w_{n})=\ell>1$, then the element $\frac{1}{m}(w_{1}/\ell,\cdots,w_{n}/\ell)$ contributes a smaller number to the right-hand-side. So we assume that $\mathrm{gcd}(w_{1},\dots,w_{n})=1$. Consider the weighted blow-up $\mathbb{C}^{n}$ with weights $(w_{1},\dots,w_{n})$. By [Kol13, pg. 104], a local chart is given by
$f:\mathbb{C}^{n}_{\mathbf{x}}/\frac{1}{w_{1}}(1,-w_{2},\dots,-w_{n})\rightarrow\mathbb{C}^{n}_{\mathbf{y}}$
with $f^{*}x_{1}=z_{1}^{w_{1}}$, $f^{*}x_{i}=z_{1}^{w_{i}}z_{i}$ for $i\geq 2$. Let $E$ denote the exceptional divisor of this weighted blowup that dominates an exceptional divisor $F$ over $\mathbb{C}^{n}/G$. Because $f^{*}(x_{1}^{r-1}dx_{1}\wedge\cdots\wedge dx_{n})=w_{1}z_{1}^{-1+(r-1)w_{1}+\sum_{i=1}^{n}w_{i}}dz_{1}\wedge\cdots\wedge dz_{n}$ and the covering $f$ is generically unramified along the exceptional divisor $\{z_{1}=0\}$, we get
$a(F,\mathbb{C}^{n}/G,(1-r)\mathcal{Y}_{0})+1=\frac{a(\{z_{1}=0\},\mathbb{C}^{n},(1-r)\{z_{1}=0\})+1}{m}=\frac{1}{m}((r-1)w_{1}+\sum_{i=1}^{n}w_{i}).$
By (13) we get the formula (14). ∎
2.3. Conley-Zehnder indices and SFT degrees
We briefly review the basics of Conley-Zehnder indices and related concepts. Given a path $\phi:[0,1]\to\mathrm{Sp}(2n)$ in the symplectic matrices, such that $\phi(0)=\mathrm{id}$ and $\phi(1)\in\mathrm{Sp}^{*}(2n):=\left\{A\in\mathrm{Sp}(2n)\left|\det(A-\mathrm{id})\neq 0\right.\right\}$, we can define the integral Conley-Zehnder index $\mu_{\mathrm{CZ}}(\phi)$, which is characterized by the following properties [Gut14, Proposition 8]:
  1. (1)
    Given any path $\psi:[0,1]\to\mathrm{Sp}(2n)$, we have $\mu_{\mathrm{CZ}}(\psi\phi\psi^{-1})=\mu_{\mathrm{CZ}}(\phi)$;
  2. (2)
    The Conley-Zehnder index is homotopy invariant for homotopies of paths from $\mathrm{id}$ to $\mathrm{Sp}^{*}(2n)$;
  3. (3)
    Given paths $\phi_{1},\phi_{2}$ in $\mathrm{Sp}(2n_{1}),\mathrm{Sp}(2n_{2})$ respectively, then the path $\phi_{1}\oplus\phi_{2}$ in $\mathrm{Sp}(2n_{1}+2n_{2})$ has Conley-Zehnder index $\mu_{\mathrm{CZ}}(\phi_{1})+\mu_{\mathrm{CZ}}(\phi_{2})$;
  4. (4)
    Given a loop $\psi:[0,1]\to\mathrm{Sp}(2n)$ with $\psi(0)=\psi(1)=\mathrm{id}$, then
    $\mu_{\mathrm{CZ}}(\phi\psi)=\mu_{\mathrm{CZ}}(\phi)+2\mu(\psi)$
    where $\mu(\psi)$ is the Maslov index of the loop $\psi$, i.e. the degree of $S^{1}\to\mathrm{Sp}(2n)\to U(n)\stackrel{{\scriptstyle\det_{\mathbb{C}}}}{{\longrightarrow}}U(1)$, here $\mathrm{Sp}(2n)\to U(n)$ is the homotopy inverse of the inclusion $U(n)\to\mathrm{Sp}(2n)$;
  5. (5)
    $(-1)^{n-\mu_{\mathrm{CZ}}(\phi)}=\mathrm{sign}\det(\mathrm{id}-\phi(1))$;
  6. (6)
    Let $S$ be a symmetric matrix such that $|S|<2\pi$, for $\phi(t)=\exp(tJS)$ where $J$ is the standard complex structure on $\mathbb{C}^{n}\simeq\mathbb{R}^{2n}$, we have $\mu_{\mathrm{CZ}}(\phi)$ is half of the signature of $S$.
Properties (2), (4) and (6) determine Conley-Zehnder index uniqueness, see [Gut14, Proposition 9].
The Conley-Zehnder indices have the following generalizations.
  1. (a)
    Conley-Zehnder index has a generalization called the Robbin-Salamon index $\mu_{\mathrm{RS}}\in\frac{1}{2}\mathbb{Z}$ defined for any path in $\mathrm{Sp}(2n)$. One important feature of $\mu_{\mathrm{RS}}$ is that it is additive w.r.t. concatenation of paths [Gut14, Property 4 of Lemma 2 6]. By property (4) above, we have $\mu_{\mathrm{RS}}(\psi)=2\mu(\psi)$ for a loop $\psi$. Moreover, we have the following normalization formula for the path in $U(1)\subset Sp(2)$:
    (15) $\mu_{RS}(\{e^{i\theta};0\leq\theta\leq T\})=\left\{\begin{array}[]{ll}\frac{T}{\pi},&\text{if }T\in 2\pi\mathbb{Z}\\ 2\lfloor\frac{T}{2\pi}\rfloor+1,&\text{if }T\not\in 2\pi\mathbb{Z}.\end{array}\right.$
  2. (b)
    $\mu_{\mathrm{CZ}}$ has a lower semi-continuous extension $\mu_{\mathrm{LCZ}}\in\mathbb{Z}$ to all paths in $\mathrm{Sp}(2n)$ starting from $\mathrm{id}$ [McL16, Definition 2.11]. By [McL16, Lemma 4.10],
    (16) $\mu_{\mathrm{LCZ}}(\phi)=\mu_{\mathrm{RS}}(\phi)-\frac{1}{2}\dim\ker(\phi(1)-\mathrm{id}).$
We now turn to the geometric setting. For a contact manifold $(M^{2n-1},\eta)$, let $\phi_{t}$ denote the flow generated by the Reeb vector field ${\mathfrak{v}}$. For a closed Reeb orbit $\gamma$, after choosing the trivialization of $\gamma^{*}(\xi:=\ker\eta)$, the linearized flow gives rise to a path in $\mathrm{Sp}(2n-2)$. By properties (2) and (4), the Conley-Zehnder index (including $\mu_{\mathrm{LCZ}},\mu_{\mathrm{RS}}$) only depends on the induced trivialization $\tau$ of $\det_{\mathbb{C}}\xi$, which will be denoted by $\mu_{\mathrm{CZ}}^{\tau}(\gamma)$ ($\mu^{\tau}_{\mathrm{LCZ}}(\gamma),\mu^{\tau}_{\mathrm{RS}}(\gamma)$) respectively. To get Conley-Zehnder indices for all Reeb orbits, we have the following:
  1. (i)
    For any Reeb orbit $\gamma$, $\mu^{\mathbb{Z}_{2}}_{\mathrm{LCZ}}(\gamma)\in{\mathbb{Z}_{2}}$ is well-defined. This is defined by $\mu_{\mathrm{LCZ}}^{\tau}(\gamma)\mod 2$ for any trivialization $\tau$ of $\det_{\mathbb{C}}\gamma^{*}\xi$. The well-definedness follows from Property (4).
  2. (ii)
    If $c_{1}(\xi)=0$, $\det_{\mathbb{C}}\xi$ can be trivialized over $M$ (the set of trivializations is parameterized by $H^{1}(M;\mathbb{Z})$). After choosing a trivialization $\tau$ of $\det_{\mathbb{C}}\xi$, we get $\mu^{\tau}_{\mathrm{LCZ}}(\gamma)$ for all Reeb orbits and it is independent of $\tau$ if $[\gamma]=0\in H_{1}(M;\mathbb{Q})$. In the case, $\mu^{\mathbb{Z}_{2}}_{\mathrm{LCZ}}(\gamma)=\mu_{\mathrm{LCZ}}(\gamma)\mod 2$.
  3. (iii)
    If $c^{\mathbb{Q}}_{1}(\xi)=0$ in $H^{2}(M;\mathbb{Q})$, we can trivialize $\det_{\mathbb{C}}\oplus^{N}\xi$ for some $N\in\mathbb{Z}_{>0}$, after choosing a trivialization $\tau$, this allows us to define $\mu_{\mathrm{LCZ}}^{\mathbb{Q},\tau}(\gamma)=\mu_{\mathrm{LCZ}}^{\tau}(\oplus^{N}\phi)/N\in\mathbb{Q}$, where $\phi$ is the linearized flow along $\gamma$. This definition is independent of $N$ and is moreover independent of $\tau$ if $[\gamma]=0\in H_{1}(M;\mathbb{Q})$ [GZ21, §3.3.2]. This rational index coincides with the integer index above when $c_{1}(\xi)=0$. In general, one should think $(\mu^{\mathbb{Q}}_{\mathrm{LCZ}},\mu^{\mathbb{Z}_{2}}_{\mathrm{LCZ}})\in\mathbb{Q}\times\mathbb{Z}_{2}$ is a pair of indices associated to an orbit.
Definition 2.12.
For a $12n-$ 1 dimensional contact manifold $(M.\xi)$ with rational first Chern class $c^{\mathbb{Q}}_{1}(\xi)=0$ and $H^{1}(M;\mathbb{Q})=0$, let $\gamma$ be a Reeb orbit for a contact form $\eta$, we define the lower SFT degree
$\mathrm{lSFT}(\gamma):=\mu^{\mathbb{Q}}_{\mathrm{LCZ}}(\gamma)+n-3.$
Under the same assumption, McLean [^2] The Robbin-Salamon index was denoted as $\mathrm{CZ}(\gamma)$ in [McL16] [McL16] defined the highest minimal index by
$\mathrm{hmi}(M):=\sup_{\eta}\inf_{\gamma}\left\{\mathrm{lSFT}(\gamma)\right\}\in\mathbb{Q}\cup\{\pm\infty\},$
where $\eta$ is a contact form and $\gamma$ is a Reeb orbit of $\eta$. When $M$ is the contact link of an isolated $\mathbb{Q}$-Gorenstein normal singularity, McLean [McL16, Theorem 1.1] proved that $\mathrm{hmi}(M)$ is twice the minimal discrepancy number if $\mathrm{hmi}(M)\geq 0$ and $\mathrm{hmi}(M)<0$ is equivalent to that the minimal discrepancy number is negative.
3. Proof of main results
In the following, we study the Fano index of the orbifold $\mathcal{Y}$ as well as the minimal discrepancy number of the cone singularity $\mathcal{C}(Y,\mathcal{L})$ using Floer cohomology from symplectic geometry to finish the proof of the main theorem.
3.1. Quasi-regular Sasaki manifolds
Return to the quasi-regular Sasaki case, the contact manifold $M$ has a contact form $\eta$ whose Reeb vector field ${\mathfrak{v}}$ generates an $S^{1}$-action and $\mathcal{Y}=M/S^{1}$ is a symplectic orbifold. The standing topological assumption in this subsection is that
$c^{\mathbb{Q}}_{1}(M)=0$
which guarantees a $\mathbb{Q}$-valued indices ($\mu_{\mathrm{LCZ}},\mu_{\mathrm{RS}}$) after choosing a trivialization of $\det_{\mathbb{C}}\oplus^{N}\xi$, and is a consequence of the cone singularity being $\mathbb{Q}$-Gorenstein. In fact, it is equivalent to the singularity being $\mathbb{Q}$-Gorenstein given that the singularity is klt by [McL16, Lemma 3.3] and [BdFFU14, Corollary 5.17]. In the special case here, we do not need to assume $H^{1}(M;\mathbb{Q})=0$, as all Reeb orbits of a conical contact form are rationally null-homologous making the index independent of the trivialization. Moreover, in the question of compactification of $\mathbb{C}^{n}$, we have $H^{2}(M;\mathbb{Z})=0$ by 2.5 guaranteeing $\mathbb{Z}$-valued indices. Let $N$ be a positive integer such that $Nc_{1}(\xi)=0$. Note that for any $x\in M$, the direct sum of $N$ copies of the linearization of the $S^{1}$-action along the orbit $S^{1}\cdot x$ gives rise to a loop in $\mathrm{Sp}((2n-2)N)$ by trivializing $\oplus^{N}\xi\subset TM$ along the orbit, where the trivialization matches with the unique (by our topological assumptions) global trivialization of $\det_{\mathbb{C}}\oplus^{N}\xi$. We use $R(M)$ to denote the Maslov index $\mu$ of this loop divided by $N$, which is independent of the point $x$ by the homotopy invariant property of the Maslov index.
Lemma 3.1.
Let $(M,\xi)$ be a contact manifold with $c_{1}^{\mathbb{Q}}(\xi)=0$ and $W$ a symplectic filling. For a rationally null-homologous Reeb orbit $\gamma$, and a bounding surface $u:S\to W$ such that $\partial u=\gamma$, using the unique trivialization of $u^{*}\det_{\mathbb{C}}TW$ over $S$, we get an integer lower Conley-Zehnder index $\mu^{u}_{\mathrm{LCZ}}(\gamma)$. Then the rational lower Conley-Zehnder index
$\mu^{\mathbb{Q}}_{\mathrm{LCZ}}(\gamma)=\mu_{\mathrm{LCZ}}^{u}(\gamma)-2\langle c_{1}(W),u_{*}[S]\rangle,$
here we view $c_{1}(W)$ as in $H^{2}(W,M;\mathbb{Q})$ as $c_{1}^{\mathbb{Q}}(\xi)=0$ and $u_{*}[S]\in H_{2}(W,M;\mathbb{Z})$. Alternatively, let $A\in H_{2}(W;\mathbb{Q})$ such that $[A]=u_{*}[S]$ in $H_{2}(W,M;\mathbb{Q})$, such $A$ exists as $u_{*}[\partial S]=0\in H_{1}(M;\mathbb{Q})$. Then the rational lower Conley-Zehnder index $\mu^{\mathbb{Q}}_{\mathrm{LCZ}}(\gamma)=\mu_{\mathrm{LCZ}}^{u}(\gamma)-2\langle c_{1}(W),A\rangle$.
Proof.
We assume $Nc_{1}(\xi)=0$, and write $\phi(t)$ as the linearized flow, applying Property (4) to $\oplus^{N}\phi(t)$ for trivializations from $u$ and the global trivialization of $\det_{\mathbb{C}}\oplus^{N}\xi$, which are differed by $2N\langle c_{1}(W),u_{*}[S]\rangle$. ∎
Under the above setup, by the $\mathbb{Q}$-Gorenstein property or $c_{1}^{\mathbb{Q}}(\xi)=0$, there is $r(\mathcal{Y},\mathcal{L})\in\mathbb{Q}$ such that $-K_{\mathcal{Y}}=r(\mathcal{Y},\mathcal{L})\mathcal{L}$. In the case of 2.5, $r(\mathcal{Y},\mathcal{L})$ is the Fano index of $\mathcal{Y}$.
Proposition 3.2.
Under the setup above, we have the identity $R(M)=r(\mathcal{Y},\mathcal{L})$.
Proof.
We can compute the Maslov index by the induced $S^{1}$-symplectic action on the symplectization $\mathbb{R}_{s}\times M$, where the symplectic tangent bundle along an orbit is the contact structure plus $\langle{\mathfrak{v}},\partial_{s}\rangle$. Along with the obvious trivialization of $\langle{\mathfrak{v}},\partial_{s}\rangle$, the linearized action along the orbit has an extra block of $2\times 2$ identity matrix compared to the contact case. In particular, the Maslov index does not change. If we compute the Maslov index using the induced trivialization from the bounding disk $D$ in the orbifold line bundle $\mathcal{L}^{-1}$, then the loop of the symplectic matrices is just a full rotation in the two coordinates in the disk $D$, hence the Maslov index is $1$. As $12R(M)-n+$ 1 is $\mu^{\mathbb{Q}}_{\mathrm{LCZ}}$ of the principle Reeb orbit (the $S^{1}$ orbit that is not multiply covered) on $M$, $R(M)$ enjoy the same property as $\mu^{\mathbb{Q}}_{\mathrm{LCZ}}$ in 3.1 without the factor $2$. Let $A$ be a closed rational chain whose intersection number with $\mathcal{Y}$ is $1$, then $[A]=[D]$ in $H_{2}(\mathcal{L}^{-1},M;\mathbb{Q})$ and $\langle c_{1}(\mathcal{L}^{-1}),A\rangle=1$. We have $R(M)=1-\langle c_{1}(\mathcal{Y})+c_{1}(\mathcal{L}^{-1}),A\rangle=r(\mathcal{Y},\mathcal{L})$. ∎
We consider the finite set $S$ of isotropy groups (including the trivial group, which is the isotropy group for a generic point) of the $S^{1}$ action on $M$. $S$ is equipped with a partial order, we say $G_{x}>G_{y}$ if $G_{y}\subset G_{x}\subset S^{1}$ is a subgroup. For $G\in S$, the quotient of the fixed point set $M^{G}/S^{1}$ gives rise to a branch locus of the quotient Kähler orbifold $\mathcal{Y}=M/\langle{\mathfrak{v}}\rangle$ giving $M$ a stratification over the partial order set $S$. For non-minimal $G\in S$, we use $G^{-}$ to denote the unique maximal element that is smaller than $G$. We formally define $G^{-}=\emptyset$ when $G$ is the minimal element of $S$. We may rescale the contact form, such that the principle orbit (the simple Reeb orbit over a non-singular point of $M/S^{1}$) has a period $1$.
Proposition 3.3.
The Reeb flow is Morse-Bott non-degenerate with the following properties:
  1. (1)
    The Morse-Bott families of Reeb orbits are parameterized by $G\in S$, $k\in G\backslash G^{-}$ and $\ell\in\mathbb{Z}_{\geq 0}$, where the family of parameterized Reeb orbit is diffeomorphic to $M^{G}$ via the starting point of the Reeb orbit and the period of the orbit is $\ell+\frac{k}{|G|}$ (we parameterize $S^{1}$ by $[0,1]$). The corresponding Morse-Bott family of unparameterized Reeb orbits is diffeomorphic to the orbifold $M^{G}/S^{1}$, which is a branch locus $\mathcal{Y}_{G}$ of the base orbifold $\mathcal{Y}$ with isotropy $G$. We will call $(G,k,\ell)$ is the signature of the Reeb orbit.
  2. (2)
    The lower Conley Zehnder index $\mu_{\mathrm{LCZ}}(G,k,\ell)$ and Robbin-Salamon index $\mu_{\mathrm{RS}}(G,k,\ell)$ of orbits with signature $(G,l,\ell)$ have the following properties:
    1. (a)
      $\mu_{\mathrm{LCZ}}(G,k,\ell)=\mu_{\mathrm{RS}}(G,k,\ell)-\dim_{\mathbb{C}}{\mathcal{Y}_{G}}$;
    2. (b)
      $\mu_{\mathrm{RS}}(\{e\},0,\ell)=2\ell R(M)$;
    3. (c)
      $\mu_{\mathrm{RS}}(G,k,\ell)=\mu_{\mathrm{RS}}(G,k,0)+2\ell R(M)$;
    4. (d)
      $\mu^{\mathbb{Z}_{2}}_{\mathrm{LCZ}}(G,k,\ell)=n-1$.
  3. (3)
    If $R(M)\geq 0$, then
    $\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\min_{G,k}\{\mu_{\mathrm{RS}}(G,k,0)-\dim_{\mathbb{C}}\mathcal{Y}_{G}+n-3\}\cup\{2R(M)-2\},$
    where $\gamma$ is a Reeb orbit of $\eta$.
Proof.
The description of the Morse-Bott family follows from that the Reeb flow is the $S^{1}$-action. $\mu_{\mathrm{LCZ}}(G,k,\ell)=\mu_{\mathrm{RS}}(G,k,\ell)-\dim_{\mathbb{C}}{\mathcal{Y}_{G}}$ follows from (16). Since the Robin-Salamon index for a loop of symplectic matrices is twice the Maslov index, we have $\mu_{\mathrm{RS}}(\{e\},e,\ell)=2\ell R(M)$. The linearized flow for $\mu_{\mathrm{RS}}(G,k,\ell)$ is the linearized flow for $\mu_{\mathrm{RS}}(G,k,0)$ concatenated with $\ell$-multiple of the loop of symplectic matrices from the $S^{1}$ action, then $\mu_{\mathrm{RS}}(G,k,\ell)=\mu_{\mathrm{RS}}(G,k,0)+2\ell R(M)$ by the concatenation property. To see 2(d), the mod $2$ lower Conley Zehnder index has two components: one from $M^{G}$ which is $\dim_{\mathbb{C}}\mathcal{Y}_{G}$ as it is a loop of symplectic matrices in $\mathrm{Sp}(\dim_{\mathbb{R}}\mathcal{Y}_{G})$ and one from the normal direction, which is $\frac{1}{2}(\dim M-\dim M^{G})=n-1-\dim_{\mathbb{C}}\mathcal{Y}_{G}\mod 2$ by property (5) of the Conley-Zehnder indices, as the endpoint/linearized return map is complex-linear. Then 2(d) follows from property (3). Since
$\mathrm{lSFT}(\gamma)=\mu_{\mathrm{RS}}(G,k,\ell)-\dim_{\mathbb{C}}\mathcal{Y}_{G}+n-3,$
When $R(M)\geq 0$, 2(b) implies that $\inf_{\gamma}\{\mathrm{lSFT}(\gamma)\}$ is attained by Reeb orbits with $\ell=0$, i.e.
$\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\min_{G,k}\{\mu_{\mathrm{RS}}(G,k,0)-\dim_{\mathbb{C}}\mathcal{Y}_{G}+n-3\}\cup\{2R(M)-2\}.$
In the following, we compute directly that $\inf_{\gamma}\mathrm{lSFT}(\gamma)$ is twice the minimal discrepancy in (14) for the conical contact form $\eta$. We will give two proofs of (17). The one given below uses axiomatic properties of the Conley-Zehnder index. Another proof is contained in Appendix A.
Proposition 3.4.
Let $o\in\mathcal{C}$ be an isolated Fano cone singularity of dimension $n$. For any conic contact form $\eta$, we have
(17) $\inf_{\gamma}\mathrm{lSFT}(\gamma)=2\mathrm{md}(o,\mathcal{C}).$
Proof.
We consider the orbifold line bundle $\mathcal{L}^{-1}$ as a symplectic filling of the link $M$. The local symplectic model of $\mathcal{L}^{-1}$ is ×$\mathbb{C}\times\mathbb{C}^{n-1}$ with cyclic action $\frac{1}{m}(w_{1},\ldots,w_{n})$, where the first factor is fiber direction and $(m,w_{1})=1$ as the link is smooth. Let $\gamma$ denote the Reeb orbit $(e^{2\pi w_{1}\sqrt{-1}t/m},0)$ for $t\in[0,1]$. We consider the trivialization $\gamma^{*}T(\mathbb{C}\times\mathbb{C}^{n-1}/\frac{1}{m}(w_{1},\ldots,w_{n}))$,
$S^{1}\times\mathbb{C}^{n}\to\gamma^{*}T(\mathbb{C}\times\mathbb{C}^{n-1}/\frac{1}{m}(w_{1},\ldots,w_{n}))$
$(t,v)\mapsto(e^{2\pi w_{1}\sqrt{-1}t/m},0)\oplus(e^{2\pi w_{1}\sqrt{-1}t/m}v_{1},\ldots,e^{2\pi w_{n}\sqrt{-1}t/m}v_{n}).$
Under such a trivialization $\tau$, the linearized Reeb flow gives rise to the following path of symplectic matrices
(18) $t\to\mathrm{diag}(1,e^{2\pi(m-w_{2})\sqrt{-1}t/m},\ldots,e^{2\pi(m-w_{n})\sqrt{-1}t/m}),\quad t\in[0,1]$
Note that although we trivialize $\gamma^{*}T(\mathbb{C}\times\mathbb{C}^{n-1}/\frac{1}{m}(w_{1},\ldots,w_{n}))$ not just $\gamma^{*}\xi$, the linearized flow in the fiber direction is identity, we may compute the SFT degree from the remaining block of the linearized flow. With such trivialization $\tau$, we have
$\mathrm{lSFT}^{\tau}(\gamma)=\#\{i|i\geq 2,w_{i}>0\}+2\#\{i|i\geq 2,w_{i}=0\}-\#\{i|i\geq 2,w_{i}=0\}+n-3=2n-4,$
where #$\#\{i|i\geq 2,w_{i}>0\}$ is from those axes with $w_{i}>0$ by Property (6) of the Conley-Zehnder index, $2\#\{i|i\geq 2,w_{i}=0\}$ from those axes with $w_{i}=0$ by Property (3) and (a), and #$-\#\{i|i\geq 2,w_{i}=0\}$ is $-\dim_{\mathbb{C}}\mathcal{Y}_{\mathbb{Z}/m}=-\frac{1}{2}\dim\ker(\phi(1)-\mathrm{id})$ in (16). Similarly, we have
$\mathrm{lSFT}^{\tau^{m}}(\gamma^{m})=2\sum_{i=2}^{n}(m-w_{i})-2,$
where $\tau^{m}$ is the induced trivialization on $\gamma^{m}$, under which the linearized flow is (18) for $t\in[0,m]$. Since $\gamma^{m}$ is $w_{1}$-th cover of a principle orbit, by 3.3, $\mathrm{lSFT}$ using a global trivialization is
$\mathrm{lSFT}(\gamma^{m})=2w_{1}r(M)-2.$
Therefore we can compute the discrepancy between the trivialization $\tau$ and the global trivialization and conclude that:
$\mathrm{lSFT}(\gamma)=\mathrm{lSFT}^{\tau}(\gamma)+\frac{\mathrm{lSFT}(\gamma^{m})-\mathrm{lSFT}^{\tau^{m}}(\gamma^{m})}{m}=\frac{2}{m}(\sum^{n}_{i=1}w_{i}+(r-1)w_{1})-2.$
Therefore $\inf_{\gamma}\mathrm{lSFT}(\gamma)$ is the minimal discrepancy in (14). ∎
3.2. Symplectic homology and its variants
Let $W$ be the Liouville filling of $M$ in $\mathbb{C}^{n}$ in 2.8. To every Liouville filling, one can associate with several symplectic invariants, namely the symplectic homology (over $\mathbb{Q}$) $SH_{*}(W;\mathbb{Q})$ as well as its variants, see e.g. [CO18]. Those invariants are always graded over $\mathbb{Z}_{2}$, they are generated by (non-degenerate) Reeb orbits/Hamiltonian orbits with grading from the Conley-Zehnder index. When $c_{1}(W)=0$, those invariants are graded by $\mathbb{Z}$, where the grading depends on a choice of the trivialization of $\det_{\mathbb{C}}TW$. If we only have $c^{\mathbb{Q}}_{1}(W)=0$, those invariants can be graded over $\mathbb{Q}$ with similar dependence on trivializations.
Our proof follows from the computation of the positive $S^{1}$-equivariant symplectic homology $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ of $W$ from two perspectives: (1) from functorial properties of the invariants, (2) direct computation with the help of the $S^{1}$-symmetry on the Sasaki link and the Morse-Bott spectral sequence [KvK16].
Remark 3.5.
There are many different conventions regarding symplectic (co)homology and its variants. Since the references cited here is not completely consistent, we point out their relations: the symplectic homology $SH_{*}$ and its variants in [BO13, CO18, KvK16] is isomorphic to the symplectic cohomology $SH^{*}$ and its variants in [Rit13, Sei08] by $SH^{*}=SH_{n-*}$, where $n$ is half of the dimension of the symplectic domain. The symplectic homology in [BO13, CO18, KvK16] is graded by Conley-Zehnder indices of the Hamiltonian orbits.
Proposition 3.6.
Let $W$ be the strongly pseudo-convex domain bounded by $M$ in $\mathbb{C}^{n}$ by 2.5, the positive $S^{1}$-equivariant symplectic cohomology of $W$ is
$SH^{+,S^{1}}_{*}(W;\mathbb{Q})=\left\{\begin{array}[]{cc}\mathbb{Q},&*=n+1+2m,m\in\mathbb{Z}_{\geq 0}\\ 0,&\text{else}\end{array}\right.$
Proof.
By gluing the Liouville filling $W$ from 2.8 and the Liouville cobordism in 2.9, we get a Liouville filling $\tilde{W}$ of the standard contact sphere. Then by a Theorem of Seidel and Smith [Sei08, Corollary 6.5], the symplectic homology of $\tilde{W}$ vanishes. We get a unital ring map $SH_{*}(\tilde{W};\mathbb{Q})\to SH_{*}(W;\mathbb{Q})$ from the Viterbo transfer [Rit13, Theorem 9.5], then we have $SH_{*}(W;\mathbb{Q})=0$. Then by the tautological long exact sequence [Rit13, Lemma 8.1]
$\ldots\to H^{*}(W;\mathbb{Q})\to SH_{n-*}(W;\mathbb{Q})\to SH^{+}_{n-*}(W;\mathbb{Q})\to H^{*+1}(W;\mathbb{Q})\to\ldots,$
we get that $SH^{+}_{*}(W;\mathbb{Q})$ is supported in degree $n+1$ with rank $1$ as $W$ is a homology ball by 2.5. By [BO13, Theorem 1.1], we have the following Gysin exact sequence,
$\ldots\to SH^{+}_{k}(W;\mathbb{Q})\to SH^{+,S^{1}}_{k}(W;\mathbb{Q})\stackrel{{\scriptstyle D}}{{\to}}SH^{+,S^{1}}_{k-2}(W)\to SH^{+}_{k-1}(W)\to\ldots.$
As a consequence, we have $D:SH^{+,S^{1}}_{k}(W;\mathbb{Q})\to SH^{+,S^{1}}_{k-2}(W)$ is an isomorphism when $k\geq n+3$ or $k\leq n$. The map $D$ is nilpotent in the sense that for every $x\in SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ there exists a $K\in\mathbb{Z}{\geq 0}$ depending on $x$, such that $D^{k}x=0$. This follows from the definition of $D$ [BO13] as it decreases the action. Therefore we have $SH^{+,S^{1}}_{k}(W;\mathbb{Q})=0$ when $k\leq n$. Then from the Gysin exact sequence, we have
$\ldots\to 0\to SH^{+}_{n+1}(W;\mathbb{Q})\to SH^{+,S^{1}}_{n+1}(W;\mathbb{Q})\stackrel{{\scriptstyle D}}{{\to}}0\to\ldots,$
and
$\ldots\to 0\to SH^{+,S^{1}}_{n+2}(W;\mathbb{Q})\stackrel{{\scriptstyle D}}{{\to}}0\to\ldots,$
That is $SH^{+,S^{1}}_{n+1}(W;\mathbb{Q})=\mathbb{Q}$ and $SH^{+,S^{1}}_{n+2}(W;\mathbb{Q})=0$. Then the claim follows from $D$ is an isomorphism when $k\geq n+3$. ∎
The next result works for more general $M$, namely those 3.3 applies. Let $W$ be a Liouville symplectic filling of $M$ such that the rational first Chern class $c^{\mathbb{Q}}_{1}(W)=0$, we have the following Morse-Bott spectral sequence [KvK16, Theorem 5.4. (13)] computing $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$.
$E_{p,q}^{1}=\bigoplus_{p=N(\ell+\frac{k}{|G|})}H_{p+q-\mu_{\mathrm{LCZ}}(G,k,\ell)}(Y_{G};\mathbb{Q})$
where $N=\mathrm{lcm}_{G\in S}(|G|)$. Here the spectral sequence is from the filtration by the period recorded by $p$ and $p+q$ is the (rational) grading on $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ by the (rational) Conley-Zehnder indices. This spectral sequence is induced by the increasing filtration by subcomplexes generated Reeb orbits with period up to $p/N$. This is a homological spectral sequence with differential $d^{r}_{p,q}:E^{r}_{p,q}\to E^{r}_{p-r,q+r-1}$. Moreover, each element from $H_{*}(\mathcal{Y}_{G};\mathbb{Q})$ also has a well-defined $\mathbb{Z}_{2}$ grading corresponding to the same spectral sequence from the period filtration but graded by the mod $2$ Conley-Zehnder index. The differential $d^{r}$ is degree $-1$ for both $\mathbb{Q}$-grading and $\mathbb{Z}_{2}$-grading. In the following picture, we spell out the $\mathbb{Q}$-grading and use different colors to mark the $\mathbb{Z}_{2}$-grading of each element.
$d^{1}$$d^{1}$$d^{2}$$p$$q$
Figure 1. A schematic picture of the first page of the spectral sequence
To present the first page of the spectral sequence, the vertical axis is $p$ and the horizontal axis is $q$, each dot represents a copy of $\mathbb{Q}$ with color indicating their $\mathbb{Z}_{2}$ grading. Each dashed block represents some $H_{*}(\mathcal{Y}_{G};\mathbb{Q})$ with the $p+q$ coordinate of the leading term is $\mu_{\mathrm{LCZ}}(G,k,\ell)$, all the leading dots have the same $\mathbb{Z}_{2}$ grading by (2d) of 3.3, where red means $n-1\mod 2$. The arrows are differentials that respect both gradings.
Proposition 3.7.
Under the assumption of 3.3 and $R(M)>0$. If we have a Liouville filling $W$ of $Y$ such that $c_{1}^{\mathbb{Q}}(W)=0$, we have
$\mathrm{hmi}(M)=\inf\{k|SH^{+,S^{1}}_{k}(W;\mathbb{Q})\neq 0\}+n-3=\inf_{\gamma}(\mathrm{lSFT}(\gamma)),$
where $\gamma$ ranges over Reeb orbits of $\eta$.
Proof.
Since $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ is an invariant independent of $\eta$ whose underlying cochain complex is supported in degrees at least $\inf_{\gamma}(\mathrm{lSFT}(\gamma))+3-n$ for any contact form $\eta$. Therefore for any contact form $\eta$, we have
$\inf_{\gamma}(\mathrm{lSFT}(\gamma))\leq\inf\{k|SH^{+,S^{1}}_{k}(W;\mathbb{Q})\neq 0\}+n-3$
where we ennumerate $\gamma$ through all Reeb orbits of $\eta$. Hence we know that
$\mathrm{hmi}(Y)\leq\inf\{k|SH^{+,S^{1}}_{k}(W;\mathbb{Q})\neq 0\}+n-3.$
Now we return to the conical contact form case. As $R(M)>0$, $\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\min_{G,k}\{\mathrm{RS}(G,k,0)-\dim_{\mathbb{C}}Y_{G}+n-3\}\cup\{R(M)-2\}$. We claim that an element in $H_{0}(\mathcal{Y}_{G};\mathbb{Q})$ with the minimum rational degree $p+q$ and the maximum $p$ will survive in the spectral sequence. For such an element $\alpha$ to be killed in the spectral sequence, we must have an element $\beta$ in $E^{1}_{p+r,q-r+1}$ that survives till the $r$th page and $d^{r}\beta=\alpha$.
[…]
Since the $\alpha$ and $\beta$ have different $\mathbb{Z}_{2}$ grading, we must have $\alpha$ red and $\beta$ blue, i.e. in the situation shown in the picture above. As a consequence, the minimal degree term $\gamma$ of the block containing $\beta$ must have strictly larger $p$ and no larger $p+q$ compared to $\alpha$, this contradicts with the choice of $\alpha$. Therefore $\alpha$ is non-trivial in $SH^{*}_{+,S^{1}}(W;\mathbb{Q})$ with the minimal grading $\inf_{\gamma}(\mathrm{lSFT}(\gamma))+3-n$. On the other hand, we have $\mathrm{hmi}(M)\geq\inf_{\gamma}(\mathrm{lSFT}(\gamma))$ by definition. Therefore we must achieve equality that
$\mathrm{hmi}(M)=\inf\{k|SH^{k}_{+,S^{1}}(W;\mathbb{Q})\neq 0\}+n-3=\inf_{\gamma}(\mathrm{lSFT}(\gamma))$
Remark 3.8.
Assume the Reeb flow of $\eta$ gives an $S^{1}$-action as above and $R(M)>0$. It is always true that $\mathrm{hmi}(M)$ is the $\inf_{\gamma}(\mathrm{lSFT}(\gamma))$, which is observed by a homology class, i.e. we do not need to assume the existence of a Liouville filling $W$. The existence of a Liouville filling allows us to use $\mathbb{Q}$-graded positive $S^{1}$-equivariant symplectic homology, which is the linearized contact homology w.r.t. the augmentation from $W$ [BO09]. In general, we can use the linearized contact homology w.r.t. an algebraic $\mathbb{Q}$-graded augmentation over $\mathbb{C}$. The Morse-Bott spectral sequence works for such linearized contact homology as well and the proof of 3.7 holds. As contact manifolds with $S^{1}$-Reeb flow are strongly fillable by [KL21], the contact homology does not vanish [LW11, Theorem 1 and Proposition 2.9]. Since the vanishing of contact homology is the only obstruction to the existence of an algebraic augmentation over an algebraically closed field, we can always use the linearized contact homology, which is expected to be a contact invariant [^3] The foundation of linearized contact homology is an ongoing project [ABZ] of Russell Avdek, Erkao Bao and the second author..
3.3. Proofs of the main theorems
Proof of Theorem 1.4.
It follows from (14), Propositions 3.4 and 3.7. ∎
Proposition 3.9.
The minimal discrepancy number of the cone singularity $\mathcal{C}(Y,\mathcal{L})$ is $n-1$.
Proof.
By Propositions 2.5 and 3.2, we have $R(M)>0$. Then Propositions 3.7 and 3.6 imply that $\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\inf\{k|SH^{+,S^{1}}_{k}(W;\mathbb{Q})\neq 0\}+n-3=2n-2$, which implies the minimal discrepancy number is $n-1$ by 1.4. ∎
Proof of 1.1.
When $\mathcal{Y}$ has no orbifold singularity, then Propositions 3.3 and 3.9 imply that $R(M)=n$. By 3.2, we have the Fano index $r(\mathcal{Y})=n$. Then the claim follows from [BM78, KO73]. ∎
Proof of Theorem 1.3.
By approximation, we can assume that there exists a sequence of rational ${\mathfrak{v}}_{k}$ that converges to ${\mathfrak{v}}$ as $k\rightarrow+\infty$. For each ${\mathfrak{v}}_{k}$, we apply Theorem 2.4 to get Kähler orbifold compactification of $\mathbb{C}^{n}$. By Proposition 2.10 we know that $\mathcal{C}$ has a Gorenstein isolated singularity at the vertex $o$.
We assume that Shokurov’s conjecture is true. Then $\mathcal{C}$ is a smooth affine variety with an effective torus action with a unique fixed attractive point $o\in\mathcal{C}$. By [KR82], we know that $\mathcal{C}$ is equivariant biholomorphic to $\mathbb{C}^{n}$ with a linear torus action. So the Reeb vector field of $(\mathcal{C},g_{0})\cong(\mathbb{C}^{n},g_{0})$ is exactly given by the imaginary part of some linear vector field $\sum_{i=1}^{n}w_{i}z_{i}\partial_{z_{i}}$ with $w_{i}>0,i=1,\dots,n$. Now if $g$ is Ricci flat, then $(\mathcal{C},g_{0})$ is also Ricci flat. By the volume minimization property of Reeb vector fields associated to Ricci flat Kähler cone metrics (see [MSY08]), the vector $(w_{1},\dots,w_{n})$ must minimize the normalized functional $\frac{(w_{1}+\cdots+w_{n})^{n}}{w_{1}\cdot\cdots\cdot w_{n}}$. This forces $w_{1}=w_{2}=\cdots=w_{n}$ and $g_{0}$ is the flat metric on $\mathbb{C}^{n}$. In other words, the metric tangent cone of $(\mathbb{C}^{n},g)$ at infinity is the flat metric on $\mathbb{C}^{n}$. By the rigidity result from [And90], we know that the Ricci flat metric $g$ on $\mathbb{C}^{n}$ itself must be flat. ∎
Appendix A An alternative proof of Proposition 3.4
We will use the same notation as our proof of the formula (14). Let $\mathcal{C}=\mathcal{C}(\mathcal{Y},\mathcal{L})$ be an isolated Fano cone singularity. Let $\mathcal{C}^{\prime}$ denote the total space of the orbifold line bundle $\pi:\mathcal{L}^{-1}\rightarrow\mathcal{Y}$ and $\mu_{1}:\mathcal{C}^{\prime}\rightarrow\mathcal{C}$ be the contraction of the zero section. Fix any point $p\in\mathcal{Y}$, by [Kol04, 25] there exist a neighborhood $U$ of $p\in\mathcal{Y}$ and $m\in\mathbb{Z}_{>0}$ and $e_{2},\dots,e_{n}\in\{0,1,\dots,m-1\}$ such that
$\pi^{-1}(U)=\mathbb{C}\times\mathbb{C}^{n-1}/\frac{1}{m}(1,e_{2},\dots,e_{n}).$
Let $\{x_{1},x_{2},\dots,x_{n}\}$ denote the coordinates on the ramified cover $\mathbb{C}^{n}$. We have the isomorphism
$\pi^{-1}(U)=\mathrm{Spec}(\mathbb{C}[x_{1}^{m},x_{2}x_{1}^{-e_{2}},\dots,x_{n}x_{1}^{-e_{n}}]).$
Assume that $-K_{\mathcal{Y}}=r\mathcal{L}$ so that $\mathcal{C}$ is $\mathbb{Q}$-Gorenstein, i.e. there exists $N\in\mathbb{Z}_{>0}$ such that $N\cdot K_{\mathcal{C}}$ is Cartier. Then we can write down a local trivializing section of $(K_{\mathcal{C}})^{\otimes N}$ as ${\sigma}_{\mathcal{C}}=\left(x_{1}^{r-1}dx_{1}\wedge dx_{2}\wedge\cdots\wedge dx_{n}\right)^{\otimes N}$ which induces a trivializing section of $((\det\xi)^{\vee})^{\otimes N}$ as ${\sigma}:=(\iota_{x_{1}\partial_{x_{1}}}\sigma_{\mathcal{C}})^{\otimes N}=(x_{1}^{r}dx_{2}\wedge\cdots\wedge dx_{n})^{\otimes N}$. Choose an orbifold Hermitian metric $h$ on the orbifold line bundle $\mathcal{L}^{-1}\rightarrow\mathcal{Y}$ such that its Chern curvature $\sqrt{-1}\partial\bar{\partial}\log h$ is an orbifold Kähler form. Locally over the uniformization chart $h$ is represented as $a(x^{\prime})|x_{1}|^{2}$ where $x^{\prime}=\{x_{2},\dots,x_{n}\}$. For any $\beta>0$, $h^{\beta}$ induces a radius function $\mathfrak{r}$ on $\mathcal{C}$ which in turn defines a Riemannian cone metric $g_{\mathcal{C}}$ on $\mathcal{C}$. Over the regular part $\mathcal{C}^{\circ}$, we can find a basis of the contact distribution $\xi$ that are compatible with the trivialization of $(\det\xi)^{\otimes N}$:
$V_{j}:=x_{1}^{e_{j}}(\partial_{x_{j}}-a^{-1}(\partial_{x_{j}}a)x_{1}\partial_{x_{1}}),\quad 2\leq j\leq n.$
Indeed, it is easy to verify that $\{V_{j}:2\leq j\leq n\}$ are basis for the orthogonal complement to the Euler vector field $x_{1}\partial_{x_{1}}$ and are invariant under the cyclic group action. Let $\nabla$ denote the Levi-Civita connection of $g_{\mathcal{C}}|_{\{\mathfrak{r}=1\}}$. Set $x_{1}=|x_{1}|e^{\sqrt{-1}\theta}$. By a straightforward calculation, we get:
$\nabla_{\partial_{\theta}}V_{j}=-(\beta+e_{j})V_{j},\quad\nabla_{\partial_{\theta}}\sigma=N(r-\beta(n-1))\sigma.$
So if we choose $\beta=\frac{r}{n-1}$ then the parallel transport trivializes $(\det\xi)^{\otimes N}$. From now on, we set $\beta$ to be of this value. Let $\gamma$ be the parametrised closed Reeb orbit: $\theta\rightarrow[(\theta,0,\dots,0)]$ with $\theta\in[0,\frac{2\pi}{m}]$. We now write down a trivialization of $\gamma^{*}\xi^{\oplus N}$ by an appropriate choice of basis that is compatible with the trivialization of $\det(\xi)^{\otimes N}$. First, we consider the Gorenstein case i.e. when $N=1$. We set
$V^{\prime}_{2}=e^{\sqrt{-1}(\beta(n-1)+\sum_{j=2}^{n}e_{j})\theta}V_{2},\qquad V^{\prime}_{j}=V_{j},\quad 3\leq j\leq n.$
Note that we have the Lie derivatives:
$\mathscr{L}_{\partial_{\theta}}V^{\prime}_{2}=\sqrt{-1}\left((n-1)\beta+\sum_{j=2}^{n}e_{j}-e_{2}\right)V^{\prime}_{2},\quad\mathscr{L}_{\partial_{\theta}}V^{\prime}_{j}=-\sqrt{-1}e_{j}V^{\prime}_{j}.$
So the linearized flow $\{\phi_{\theta}:0\leq\theta\leq 2\pi/m\}$ is represented by the diagonal matrix:
$(e^{\sqrt{-1}((n-1)\beta+\sum_{j=2}^{n}e_{j})\theta},1,\dots,1)\cdot(e^{-\sqrt{-1}e_{2}\theta},\dots,e^{-\sqrt{-1}e_{n}\theta}).$
Since $(n-1)\beta=r$, its CZ index is equal to
$2\frac{1}{m}(r+\sum_{j=2}^{n}e_{j})-\sharp\{j:e_{j}>0\}=2\frac{1}{m}(rw_{1}+\sum_{j=2}^{n}w_{j})-(n-1)+\frac{1}{2}\dim\mathrm{ker}(\phi_{2\pi/m}-\mathrm{Id})$
where $w_{1}=1,w_{j}=e_{j},2\leq j\leq n$. Similarly for $1\leq k<m$, the CZ index of $\gamma^{k}$ is calculated by using the path $\{\phi_{\theta},0\leq\theta\leq 2\pi\frac{k}{m}\}$ (see (15)):
$2\frac{k}{m}\left(r+\sum_{j=2}^{n}e_{j}\right)-\sum_{j:e_{j}>0}\left(2\left\lfloor\frac{e_{j}k}{m}\right\rfloor+1\right)=2\frac{1}{m}\left(kr+\sum_{j:e_{j}>0}(ke_{j}\;\mathrm{mod}\,m)\right)-\sharp\{j:e_{j}>0\}$
$=$ $2\frac{1}{m}(rw_{1}+\sum_{j=2}^{n}w_{j})-(n-1)+\frac{1}{2}\dim\mathrm{ker}(\phi_{2\pi k/m}-\mathrm{Id})$
where we used the identity $(w_{1},w_{2},\dots,w_{n})=(k,ke_{2}\,\mathrm{mod}\,m,\dots,ke_{n}\,\mathrm{mod}\,m)$. When $k=m$, $\mathrm{lSFT}(\gamma^{m})$ is $22r-$ 2. So we indeed get:
(19) $\inf_{1\leq k\leq m}\mathrm{lSFT}(\gamma^{k})=2\min\left\{r,\frac{1}{m}(rw_{1}(g)+\sum_{j=2}^{n}w_{j}(g)),g\in G\right\}-2.$
When $m>k$, $\mathrm{lSFT}$ can be computed similarly but will be strictly larger than (19) as $r>0$. Finally for the case when $N\geq 2$, we can trivialize $\gamma^{*}\xi^{\oplus N}$ by using the following basis vectors:
$V^{\prime(p)}_{j}=\left\{\begin{array}[]{ll}e^{\sqrt{-1}N(\beta(n-1)+\sum_{j=2}^{n}e_{j})\theta}V_{2},&\text{if}\;p=1,j=2\\ V_{j},&\text{otherwise i.e. if }p=1,3\leq j\leq n\text{\; or\; }2\leq p\leq N,2\leq j\leq n.\end{array}\right.$
Then a similar calculation gives the same expression as (19).